OpenStudy (anonymous):

*************medals and fans************Cormac is making 29 phone calls each week as part of a fundraiser. Last week, he made 12 phone calls on weekdays. Which equation could be used to find x, the number of phone calls he made last week on weekend days?

OpenStudy (anonymous):

29x = 12 x + 29 = 12 12 + x = 29 x - 29 = 12

OpenStudy (anonymous):

29-12=x

OpenStudy (anonymous):

sooo that would be what?

OpenStudy (anonymous):

the last problem

OpenStudy (anonymous):

Your wrong no medal for u sorry MY HUSBAND WILL HUNT U DOWN HAAAAAAA